# Standard Language Ideology in AI-Generated Language

### v1

@@ 4.1. Standard AI-generated language ideology
This is further reinforced by AI-generated language outputs being of lower quality for inputs in marginalized varieties compared to inputs in “standard” varieties. Given this, native speakers of minoritized languages may be incentivized to provide inputs in “standard” varieties even if they are less fluent in them, a form of digital code switching. If users prompt outputs to be in the particular minoritized language variety of the input, outputs can disproportionately carry stereotypes, produce demeaning content, facilitate appropriation, or support manipulation. Taken together, these harms can impact access to resources or opportunities when generative AI tools are used as gateways or checkpoints (e.g., chatbots for scheduling healthcare appointments or educational tutors) and as they become increasingly integrated into daily life. It illustrates the powerful—yet subtle—ways in which Western hegemony, and particularly American hegemony, manifests in and through emerging AI technologies.
@@ 4.2. No clear way for popular language models to “win”
4.3. Moving towards emancipatory outcomes
Instead of asking what the “appropriate” way is for generative AI models to behave, perhaps a more important question is: how might generative AI models be developed to support more emancipatory outcomes, and what do emancipatory outcomes from generative AI look like?
@@ 4.3. Moving towards emancipatory outcomes
Perhaps most simply, the fact that language models perpetuate discrimination on the basis of language variety means that evaluation of language models for potential harms against minoritized groups (e.g. “toxicity” evaluations), which often focus on common demographic attributes such as race or gender, should expand to capture discrimination based on the user’s language variety. Measurement of harms due to standard language ideology is particularly important because language use often serves as a proxy for other harms–a way that racism, sexism, or xenophobia may fly under the radar.
@@
Issues surrounding appropriate responses to minoritized varieties, and when/if responding in the variety is appropriate, stem in part from the fact that these models are developed by organizations that are not composed of members of these speech communities. Instead, the development of language models by speaker communities, for speaker communities, helps to avoid this top-down imposition of linguistic norms (see, for example, (mahelona2023openai, ; nekoto-etal-2020-participatory, ), and the broader discussion of indigenous data sovereignty by (kukutai2016indigenous, )). However, given the dominance of a small set of LLMs, such as ChatGPT, there is also an onus for developers of these widely used models to improve model quality for speakers of different language varieties. Despite the difficulty of implementing participatory practices in model design, setting “ambitious, yet reasonable goals” that help to move from one of the “extreme poles of transactional consulting and transformative empowerment” to the other (delgado2023, ) could help to improve power dynamics involving speakers of minoritized varieties. This includes innovating around participatory processes that truly center the agency of minoritized people in the design, development, and ownership necessary to move towards more emancipatory digital futures.
In answering these open questions and advancing opportunities, it is important to consider what language data these models are learning from, how they can better learn from minoritized language varieties in non-extractive ways, and how marginalized communities can be enabled as central actors in the development and ownership of generative AI tools.
@@ 5. Limitations and Ethical Considerations
This paper highlights the ways in which standard language ideology is present in AI-generated language, with implications for minoritized communities and society more broadly. While we believe that our taxonomy on open problems is comprehensive and we consider a range of issues, the list and associated implications are likely not exhaustive. A potential consequence of overlooking some work is narrowing the range of perspectives considered in future research based on the avenues we mention and the structure of our taxonomy. Nonetheless, it represents a first step in the development of a complete taxonomy and introduces critical questions to inform forward paths. Our discussion also centers largely on standard language ideology in English; potential differences in the manifestation of standard language ideology in other languages is an important area for future work.

### v2

@@ 4.1. Standard AI-generated language ideology
This is further reinforced by AI-generated language outputs being of lower quality for inputs in marginalized varieties compared to inputs in “standard” varieties. Given this, native speakers of minoritized languages may be incentivized to provide inputs in “standard” varieties even if they are less fluent in them, a form of digital code switching. If users prompt outputs to be in the particular minoritized language variety of the input, outputs can disproportionately carry stereotypes, produce demeaning content, facilitate appropriation, or support manipulation. Taken together, these harms can impact access to resources or opportunities when generative AI tools are used as gateways or checkpoints (e.g., chatbots for scheduling healthcare appointments or educational tutors) and as they become increasingly integrated into daily life. It illustrates the powerful, yet subtle, ways in which Western hegemony, and particularly American hegemony, manifests in and through emerging AI technologies.
@@ 4.2. No clear way for popular language models to “win”
5. Recommendations: Moving towards emancipatory outcomes
Instead of asking what the “appropriate” way is for generative AI models to behave, we proposal a more foundational question: how might the recognition of standard language ideology guide the development of generative AI models towards more emancipatory and socially just outcomes? This reframing is crucial because there is no universally ”correct” way for models to respond across all contexts (e.g., there is not a clear answer around what the default production of models should be), and a narrow focus on performance can obscure deeper social and political stakes that standard language ideology calls us to examine. Therefore, rather than aligning our recommendations to each individual open problem, we instead take a higher-level perspective to focus our recommendations on addressing the underlying power dynamics that shape AI systems, including questions of ownership, control, and community agency. We present three concrete actions for AI researchers, funders, and practitioners.
@@ 5. Recommendations: Moving towards emancipatory outcomes
1. Expand evaluation metrics to include language variety harms. First, and most simply, the fact that language models perpetuate discrimination on the basis of language variety means that evaluation of language models for potential harms (e.g. “toxicity” evaluations) should expand to capture discrimination based on the user’s language variety. Measurement of harms due to standard language ideology is particularly important because language use often serves as a proxy for other harms–a way that racism, sexism, or xenophobia may fly under the radar.
We therefore recommend that AI developers and researchers do the following: (a) Incorporate language variety as a specific category in model evaluation frameworks alongside race, gender, and other demographic attributes; (b) develop benchmarks and metrics that assess whether models perpetuate harms (e.g., erasure of linguistic features, stereotyping, and condescension) against different minoritized language varieties; and (c) publicly share evaluation results to ensure transparency and accountability, while raising awareness about the risks of reinforcing linguistic discrimination and hierarchies through AI generated language. By incorporating language variety as a category for evaluating performance and ”toxicity” and being transparent about results, we can better capture and understand the implications of AI technologies on marginalized language communities globally.
@@
2. Support speakers of different language communities in creating their own datasets and models. Issues surrounding appropriate responses to minoritized varieties, and when/if responding in the variety is appropriate, stem in part from the fact that these models are developed by organizations that are not composed of members of these speech communities. Instead, the development of language models by speaker communities, for speaker communities, helps to avoid this top-down imposition of linguistic norms (see, for example, (mahelona2023openai, ; nekoto-etal-2020-participatory, ), and the broader discussion of indigenous data sovereignty by (kukutai2016indigenous, )).
There are various examples of communities whereby speakers of minoritized language varieties are collecting data and developing language models. One example is Masakhane (orife2020masakhane, ), a grassroots community focused on NLP research “in African languages, for Africans, by Africans.” With over one thousand participants in their community across 30 African countries, they have developed translation results for approximately 40 African languages and supported open-source datasets and subsequent machine translation models to address under-representation of African languages in technology. Another example is AI4Bharat (kunchukuttan2020ai4bharat, ), a research lab at IIT Madras that ”is dedicated to advancing AI technology for Indian languages through open-source contributions.” The organization seeks to build language models and datasets for various Indian languages across India’s diverse linguistic landscape. We recommend additional support for these types of initiatives in which speakers of different communities are collecting data and developing models for their communities. In providing support to such initiatives, ensuring data collection is not extractive and providing funding for data collection efforts and people whose data is collected remains important.
There are some cases that go further by ensuring that people of their community have a say in how that data is used. Te Hiku Media – a Māori-owned and operated nonprofit based in New Zealand – focuses on preserving and promoting the Māori language (te reo Māori) and culture (james2020developing, ). In 2018, they launched a competition in which Māori speakers across New Zealand recorded over 300 hours of annotated audio. With this, they built AI tools for automatic speech recognition and pronunciation of te reo Māori. Te Hiku Media created various AI technologies and tools for Māori (e.g., automatic speech recognition, auto tagging, real time language feedback). They also established new data protection protocols and worked with Māori-led organizations to develop AI tools that support the community. Simultaneously, they worked with other Indigenous communities to leverage the Te Hiku language database to train language models for other Indigenous groups. The diversity of methods adopted by these organizations–cross-national collaboration, open-source contributions, and community-centered competitions–illustrates how unconventional methods of research can help to center the needs of non-“standard” speech communities. In 2024, a data sovereignty agreement between Microsoft and Te Tumu Paeroa (the office of the Māori trustee) was established, in which Te Tumu Paeroa, acting as the anchor tenant for Microsoft’s New Zealand-based data centers, ensured that Māori data would be stored locally and governed according to a Māori data sovereignty framework. This agreement reflects an institutional recognition and honoring of Indigenous data governance that supports community ownership.
We recommend that multilateral institutions, funders, and policymakers prioritize meaningful support—financial, technical, and infrastructural—for language communities’ own dataset and AI development efforts. Community-led development helps prevent the top-down imposition of linguistic norms and enables more contextually appropriate, culturally-informed systems. As the approaches we highlight demonstrate, these types of efforts are already underway and can be further enabled with targeted support. Crucially, supporting language communities also means respecting their preferences and boundaries. In some cases, communities may choose to practice refusal—opting not to share or use their data in the development of AI technologies by outside actors—or may prioritize data sovereignty above participation. Supporting diverse language communities means supporting these choices, including if and when they resist mainstream technological integration.
3. Innovate on and implement participatory processes that center marginalized communities, while rethinking language data practices to avoid extractive approaches. Given the dominance of a small set of LLMs, such as ChatGPT, there is also an onus for developers of these widely used models to improve model quality for speakers of different language varieties (not only rely on those communities to address linguistic data and performance gaps). In this case, we recommend that AI researchers and practitioners: (a) Engage minoritized language communities as equal partners throughout the AI development lifecycle, from data collection and model training to deployment and evaluation (b) support community ownership by developing language models by and for speaker communities that reflect the values, norms and linguistic practices of communities they serve; and (c) implement participatory governance structures that give communities the power to make decisions about how their language data is used and tools are deployed.
An example here is the partnership between OpenAI and the Icelandic government. While most of Iceland’s citizens speak English or another second language, its integration with the United States and Europe has put the country’s native tongue, Icelandic, at risk. With concerns that Icelandic cannot remain the default language in the face of digitalization, the Icelandic government partnered with OpenAI to use GPT-4 in the preservation effort of the Icelandic language (iceland, ). As part of this initiative, $\eth$eind ehf, a language technology company, assembled a team of 40 volunteers to train GPT-4 on proper Icelandic grammar and cultural knowledge. In this case, AI practitioners (with the help of the particular community) recognized that the model was not serving the language community and sought to better learn from the speakers to improve model performance for the language variety. However, OpenAI is benefiting from the engagement of the particular linguistic community and it is not clear to what extent (if at all) the linguistic community is sharing in the financial benefits over time. This initiative could go further by ensuring that Icelandic speakers are central actors in the development and management of generative AI tools that use and leverage their language and the work they provided to improve the models.
Despite the difficulty of implementing participatory practices in model design, setting “ambitious, yet reasonable goals” that help to move from one of the “extreme poles of transactional consulting and transformative empowerment” to the other (delgado2023, ) could help to improve power dynamics involving speakers of minoritized varieties. In addition, participatory AI practices include compensating them fairly for their contributions and respecting indigenous data sovereignty principles. While participatory processes may require additional time and resources, they are critical to create AI systems that empower rather than further marginalize communities. AI researchers and practitioners must rethink their approach to language data by recognizing what language data these models are (and are not) learning from, how they can better learn from minoritized language varieties in non-extractive ways, and how marginalized communities can be enabled as central actors in the development and ownership of generative AI tools.
We recommend that AI researchers and practitioners: (a) Audit existing language datasets to identify gaps in representation for different language varieties and be transparent about linguistic data representation in models and about what languages the models work better (and worse) for; (b) incorporate diverse language into training data, ideally in ways that build from recommendation 2; and (c) collaboratively develop mechanisms to compensate and credit communities whose language data is used in model training. As researchers strive to move from merely consulting language communities to empowering them regarding the use and management of their data (delgado2023, ), the examples we describe in each of these directions can serve as case studies for future collaboration.
These actions require a mindset shift towards prioritizing equity considerations and valuing linguistic diversity as a fundamental aspect to AI development, as well as securing additional resources. Although these actions may increase development costs and timelines, they are critical to addressing the harms perpetuated by current language models and creating more emancipatory technological futures.
6. Conclusion
In this paper we identified how standard language ideology is generated and perpetuated by LLMs that are powering some of the largest generative AI applications used today, resulting in “standard” AI-generated language in which SAE is the linguistic default and holds the linguistic bias that SAE is the most “appropriate” language. Our taxonomy of open problems regarding standard language ideology in AI-generated language captured the implications for different language communities. While tensions remain, including questions around what desirable system behavior looks like, we argue that the focus should instead be on moving towards emancipatory outcomes. We present three actions for AI researchers, practitioners, and funders: (1) expand evaluation metrics to include language variety harms; (2) support speakers of different language communities in creating their own datasets and models; and (3) innovate on and implement participatory processes that center marginalized communities, while rethinking language data practices to avoid extractive approaches.
@@ 7. Limitations and Ethical Considerations
This paper highlights the ways in which standard language ideology is present in AI-generated language, with implications for minoritized communities and society more broadly. While we believe that our taxonomy on open problems considers a range of issues, the list and associated implications are likely not exhaustive. A potential consequence of overlooking some work is narrowing the range of perspectives considered in future research based on the avenues we mention and the structure of our taxonomy. Nonetheless, it represents a first step in the development of a complete taxonomy and introduces critical questions to inform forward paths. Our discussion also centers largely on standard language ideology in English; the manifestation of standard language ideology in other languages is an important area for future work.